At what temperature will 4.15g of CO2 exert a pressure of 855 mmHg at a volume of 650 ml? (make sure to turn grams into moles)


1
Expert's answer
2022-05-01T18:19:04-0400

Solution:

The molar mass of CO2 is 44.01 g/mol

Therefore,

Moles of CO2 = (4.15 g CO2) × (1 mol CO2 / 44.01 CO2) = 0.0943 mol CO2


Convert mmHg to atm:

1 atm = 760 mmHg

Therefore,

P = (855 mmHg) × (1 atm / 760 mmHg) = 1.125 atm


Convert mL to L:

1 L = 1000 mL

Therefore,

V = (650 mL) × (1 L / 1000 mL) = 0.65 L


R = gas constant = 0.082 L atm K−1 mol−1


Ideal gas law can be used to calculate the temperature:

PV = nRT

T = PV / nR

T = (1.125 atm × 0.65 L) / (0.0943 mol × 0.082 L atm K−1 mol−1) = 94.57 K

T = 94.57 K


Answer: at 94.57 K